OpenStudy (anonymous):

20 + n = 37 * this is an equation•

jimthompson5910 (jim_thompson5910):

20 + n = 37 20-20 + n = 37-20 n = 17

jimthompson5910 (jim_thompson5910):

You're basically asking "20 plus what number is 37?" The answer is "20 plus 17 is 37"